# LSEAttention is All You Need for Time Series Forecasting

### v2

@@ 3.4 Ablation Study
In this section, I present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework. By systematically removing or modifying key elements, such as the GeLU and the PReLU, I isolate their contributions to the overall performance of the model. For consistency, I fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
Dataset Pred. Length With PReLU With ReLU
ETTh2 96 0.290 0.292
ETTm1 96 0.300 0.305
ETTm2 96 0.176 0.191
Exchange Rate 96 0.098 0.115
Table 3: Performance of LATST with different activation functions
Preliminary results from the ablation study, as shown in Tab. 3, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets. This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
To further investigate the contribution of the Log-Sum-Exp trick, I conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
Dataset Pred. Length Without GeLU Transformer
ETTh2 96 0.312 0.396
ETTm1 96 0.333 0.384
ETTm2 96 0.183 0.200
Exchange Rate 96 0.105 0.292
Table 4: Performance of LATST without GeLU compared to Transformer
As illustrated in Tab. 4, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration. This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function. The improvement achieved through the Log-Sum-Exp trick is attributed to its ability to stabilize gradient computations during training, thereby facilitating the learning of more complex patterns.
These findings not only elucidate the effectiveness of various components within the LATST architecture but also provide valuable insights for future enhancements. I aim to explore optimal configurations of these modules to further refine the model’s performance and expand its applicability to more complex time series forecasting challenges.
@@ 4 Conclusion and Future Work
In this paper, I present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting. Our empirical results confirm the effectiveness and robustness of the proposed LATST approach across a variety of forecasting contexts. Although I do not exhaustively explore all potential causes of entropy collapse, I introduce a novel module specifically designed to address this issue.
@@
In summary, I believe that our work not only addresses critical challenges associated with the entropy collapse phenomenon but also sets the stage for future advancements in time series forecasting methodologies.

### v3

@@ 3.4 Ablation Study
In this section, we present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework. By systematically removing or modifying key elements, such as the GeLU and the PReLU, their contributions are isolated to the overall performance of the model. For consistency, we fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
Table 5: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
Dataset Without GeLU Transformer
ETTh2 0.312 0.396
ETTm1 0.333 0.384
ETTm2 0.183 0.200
Exchange Rate 0.105 0.292
Preliminary results from the ablation study, as shown in LABEL:tab:ab_relu, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets. This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
To further investigate the contribution of the Log-Sum-Exp trick, we conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
As illustrated in LABEL:tab:ab_gelu, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration. This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function. The improvement achieved through the Log-Sum-Exp trick is attributed to its ability to stabilize gradient computations during training, thereby facilitating the learning of more complex patterns.
These findings not only elucidate the effectiveness of various components within the LATST architecture, but also provide valuable insights for future enhancements. We aim to explore optimal configurations of these modules to further refine the model’s performance and expand its applicability to more complex time series forecasting challenges.
@@ 4 Conclusion and Future Work
In this paper, we present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting. Our empirical results confirm the effectiveness and robustness of the proposed LATST approach inty of forecasting contexts. Although we do not exhaustively explore all potential causes of entropy collapse, we introduce a combination of novel modules specifically combined to address this issue.
@@
In summary, we believe that our work not only addresses critical challenges associated with the entropy collapse phenomenon but also sets the stage for future advancements in time series forecasting methodologies.